OpenStudy (anonymous):

solve the absolute value equation or indicate that the equation has no solution IX-2I - 9 my answer sheet gives me {-7,11} I have gotten opposite value signs where did I go wrong?

OpenStudy (michele_laino):

If I apply the definition of absolute value, then your equation is equivalent to those 2 equations below: \[\Large \begin{gathered} x - 2 = 9 \hfill \\ \hfill \\ - x + 2 = 9 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

sorry I walked off

OpenStudy (anonymous):

thanks this way I get the right value signs